The labelling on a bottle of H2O2 solution. Is 20 volume the the concentration of H2O2 in percentage srength will be

Grade:11

1 Answers

Arun
5 years ago
This means that 1 volume of H2O2 will give off 20 volumes of O2 at 20C (298.15K) and 1.0 atm 

So, 1 liter of solution can produce 20 liters of oxygen under these conditions. Let's assume oxygen is a perfect gas, then from 

PV=nRT where P = 1 atm and V = 20 L, you can calculate n, the number of moles of O2 

n = PV/RT = (1)(20) / (0.0821 X 298.15) 

n = 0.818 moles of O2 

Knowing hydrogen peroxide decomposes as 2H2O2 ---> 2H2O + O2 

and 1 liter of your solution will generate 0.818 moles of O2, there must be 1.636 moles of H2O2 present in 1 liter. 

Molar mass of H2O2 = 34 g/mol, so 

1.636 mol/L of H2O2 = 55.62 grams per liter. 

density of H2O2 is 1.463 g/mL so 55.62 grams is about 38.02 mL 

%vol = 0.038L / 1.0 L X 100% = 3.8% 

For simplicity, you have a 4% solution of hydrogen peroxide.